OpenStudy (anonymous):

7 5/7 + 2 1/2.. steps please.. thank you for helping

OpenStudy (anonymous):

add the two integers 7 + 2 = 9 5/7 + 1/2 lcm = 14 10/14 + 7/14 = 17/14 = 1 3/14 total is 9 + 1 3/14 = 10 3/14

OpenStudy (anonymous):

5/4 ? - isnt it 5/2 ?